OpenStudy (anonymous):

MEDAL AND FAN!!! The population of Glendale this year is 3,200 more than half of last year’s population. The population this year is 26,300. Which equation could be solved to find last year’s population?

OpenStudy (anonymous):

Here are the answer choices. A. 1/2x + 3,200=26,300 B. 2x – 3,200 = 26,300 C. 1/2x - 3,200 = 26,300 D. x + 1/2x + 3,00 = 26,300

OpenStudy (tkhunny):

You tell me. Which expression indicates what?

OpenStudy (anonymous):

That's what I need help figuring out. If someone can run me through the correct process.

OpenStudy (tkhunny):

The correct process is thinking about it. If x is last year's population, What does ½x mean? What does 2x mean? What does x + ½x mean?

OpenStudy (anonymous):

1/2x= half of last years population? 2x= twice the amount of last years population? x + 1/2x= last years population plus half last years population?

OpenStudy (tkhunny):

Knowing that, can we eliminate choices?

OpenStudy (anonymous):

Yes we can eliminate A, B and C unless I'm mistaken. Also the 3,00 in D is suppost to be 3,200, but I made a typo.

OpenStudy (tkhunny):

Why are we eliminating A or C? Do they, or do they not refer to ½x?

OpenStudy (anonymous):

Sorry I got confused lol. Let's eliminate B. What next?

OpenStudy (anonymous):

I'm beginning to think it's A.

OpenStudy (tkhunny):

D starts with x and then add ½x - That's no good. C starts with ½x, but subtracts 3200. -- That's no good. B starts with 2x -- That's no good. A starts with ½x and adds 3200. That's exactly what is needed. Don't begin to think... Prove it!

OpenStudy (anonymous):

Well thanks, so much.